10 tube lights of 20 watt each are used in house for 5 hours daily calculate the bill of one month for the electricity with the rate of rupees 7.30 per unit.

Grade:10

2 Answers

Arun
4 years ago
Dear student
 
Total units for 1 bulb (kwh) = watt * hour * day /1000 = 20*5*30/1000 = 3
hence for 10 bulbs = 30 units.
 
hence total bill = 30*7.5 = 225
Vikas TU
4 years ago
Dear arun , 
Have some sense , It is for one month 
1 BULB=20 WATT
10 BULB=200*10
             =20000 WATT OR 20KILOWATT
TIME=5HOURS
​ENERGY=POWER *TIME
                 20KW * 5 HRs
               =100KWH OR 100 UNITS
1 DAY ENERGY= 100 UNITS
30 DAYS ENERGY= 100 *30 = 3000 unit 
1UNIT=7.30 RUPEES
3000 UNIT = 21900 RUPEES
HOPE THIS HELP